# CAVE: Controllable Authorship Verification Explanations

### v1

@@ Abstract
Authorship Verification (Av) (do two documents have the same author?) is essential for many sensitive real-life applications. Av is often used in proprietary domains that require a private, offline model, making SOTA online models like ChatGPT undesirable. Other SOTA systems use methods, e.g. Siamese Networks, that are uninterpretable, and hence cannot be trusted in high-stakes applications. In this work, we take the first step to address the above challenges with our model Cave (Controllable Authorship Verification Explanations): Cave generates free-text Av explanations that are controlled to be 1) structured (can be decomposed into sub-explanations with respect to relevant linguistic features), and 2) easily verified for explanation-label consistency (via intermediate labels in sub-explanations). In this work, we train a LLaMa-3-8B as Cave; since there are no human-written corpora for Av explanations, we sample silver-standard explanations from GPT-4-Turbo and distill them into a pre-trained LLaMa-3-8B. Results on three difficult Av datasets IMDb62, Blog-Auth, and FanFiction show that Cave generates high quality explanations (as measured by automatic and human evaluation) as well as competitive task accuracies111github.com/INK-USC/Controllable-AV-Explanations .
@@ 1 Introduction
Figure 1: Our model, Cave generates structured free-text explanations that can be easily and automatically verified for consistency.
@@
Authorship Verification (Av) Koppel and Winter (2014) is the NLP task of determining if two input documents were written by the same author. Av is used for tasks such as plagiarism detection, forensic analysis (often in support of law enforcement), analysis of the spread of misinformation. Given the sensitive nature of this task, it is imperative to develop Av methods that are interpretable, accessible and secure. Interpretability ensures that the system’s decisions can be trusted by user. Accessibility ensures easier downstream usability; for example, a well-structured free-text explanation with clearly defined topic points is more easily understood than unstructured explanations such as long lists of low-level features. Finally, security is necessary to safely use the Av system for proprietary data; a locally-hosted model on the user’s server is more secure than an API call to an online model, which could potentially store the user’s sensitive data. Av is also an extremely difficult task for humans, in contrast to other NLP tasks such as machine translation; this further necessitates the need for automatic Av systems that accessible to the user, accurate and explainable.
@@
Early works performed Av by comparing hand-crafted features such as n-grams van Dam (2013), POS tags Moreau and Vogel (2013), LIWC features Uchendu et al. (2020); while these are explainable, they are hard to scale. Subsequent works employed neural architectures to improve scalability and accuracy, such as the usage of Siamese networks Araujo-Pino et al. (2020); Najafi and Tavan (2022); Boenninghoff et al. (2019) to assign authorship via distance between document embeddings. These methods, however, have little or no interpretability. This poses a huge problem: one cannot, for instance, make legal assertions about the authorship of a set of questionable documents solely on the basis of “a system said so” - decision makers (judges, university officials, intelligence analysts, etc.) need to know why.
@@
Recent years have seen the surge of large language models and their self-rationalization capabilities. This leads to works such as Hung et al. (2023); Huang et al. (2024) that generate free-text rationales [^2] We use the term explanation and rationale interchangeably in this work. for the task of Av from models such as GPT-4 OpenAI (2024b) in a zero/few-shot manner. However, these methods face various issues such as unreliability of online models (varying results/details across multiple runs, deprecation of existing APIs), inconsistency of rationales with respect to predicted label Hung et al. (2023) (refer to Appendix B for examples), and high expenses.
In this work, we train a high-quality, in-house model Cave (Controllable Authorship Verification Explanations) (refer Figures 1, 2) to generate structured and consistent explanations for the task of Av. Cave’s free-text rationales include linguistic and stylistic analyses (such as writing style, expressions/idioms used) deemed necessary for Av Hung et al. (2023); Boenninghoff et al. (2019). The free-text rationales have an overall structure that enables the user to easily decompose them into constituent linguistic analyses; this makes the rationales more accessible and understandable to the user. Finally, the rationales are consistent, i.e., users can easily verify whether the sub-parts of the explanation (each of which has an intermediate task label) and the predicted label are consistent with each other.
We finetune a pretrained LLaMa-3-8B  [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B Meta-AI (2024) as Cave. Since there are no human-written corpora for the task of Av explanations, we follow model distillation strategies used by prior works West et al. (2022); Li et al. (2023a, 2024), wherein we generate silver-standard training data from a large teacher model GPT-4-Turbo and use that to train the smaller student model LLaMa-3-8B. Further, as discussed in Ramnath et al. (2023) it is important to create a balance between accuracy and quality of rationales during distillation to create the same balance in the student model; we follow Ramnath et al. (2023); Li et al. (2023a) and filter the silver standard data according to our rationale metrics (Section 2.2) before using them for distillation.
We first describe our pipeline for Cave in Section 2. We provide experiments and results in Section 3 and provide further discussions in Section 4. Our experiments with three difficult Av datasets show that Cave leads to competitive task performances as well as high quality of rationales.
@@ 2 Our Method
Our pipeline consists of three parts: (1) generating silver rationales from GPT-4-Turbo OpenAI (2024a) via a novel prompt, (2) filtering them according to our criteria of accessibility, consistency and accuracy, and (3) distilling these into a small language model to get our final model Cave. Figure 2 shows a visual demonstration of the pipeline.
@@ 2.1 Cave’s prompt
We begin with prior work PromptAV Hung et al. (2023) as a baseline. The PromptAV approach starts with a set of linguistic features relevant to the Av task Boenninghoff et al. (2019), e.g. writing style, sentence structure, tone of writing, punctuation/capitalization/abbreviation style, syntax/expressions/idioms used, etc. It incorporates these features in a prompt, asking an LLM to generate 1) an explanation that considers these features, and 2) a score from 00 to $1$ that shows the model’s confidence that the two documents were written by the same author. In this work, we extend PromptAV to improve the controllability of the explanation’s structure, and to enforce consistency between the intermediate rationale steps, the predicted score and the predicted task label.
We show Cave’s prompt and a sample output in Tables 2 and 2: we use a JSON-type output format where the linguistic features are keys; for each feature, the model has to provide an explanation and an intermediate output (YES/NO/MAYBE) that denotes whether this feature indicates shared authorship. Finally, the model has to provide a score from 00 to $1$ which indicates its confidence that the documents have the same author, and a YES/NO task label output. The output’s JSON format enables us to easily parse the explanation into its different sub-steps in an automatic manner; given the complex nature of Av, this format is necessary for the downstream accessibility of the explanation, and its automatic evaluation. We use Cave’s prompt to generate silver rationales that are filtered and used to train/distill a smaller model.
2.2 Data Filtering
We use three criteria to filter the silver rationales before we use them to train the smaller LM.
JSON format correctness and completeness.
We filter out any silver rationales whose format is incorrect or which are missing any of the target linguistic features.
Task Accuracy.
We retain only those silver rationales whose output label is correct. That is, the label is YES if the authors are indeed the same, or NO otherwise.
@@ 2.2 Data Filtering
Rationales which inconsistent with their predicted label Lyu et al. (2023) are unusable in downstream applications. Prior works check for consistency either by human analysis Chen et al. (2023b) or by training models Hase et al. (2020); Chen et al. (2023a); Wiegreffe et al. (2021) that can do the same. In Cave’s output format, the presence of intermediate and final labels allows us to automatically verify rationale-label consistency. We define a Consistency metric using two binary $0/1$ variables Cs-1 and Cs-2:
  1. 1.
    Cs-1: Is the predicted ‘final score’ (which indicates the model’s confidence in whether the two documents have the same author) consistent with the predicted label? If the predicted label is ‘YES’, is the final score $\geq 0.5$? Similarly, if the predicted label is ‘NO’, is the final score $\leq 0.5$?
  2. 2.
    Cs-2: Are the intermediate labels faithful to the final label? Since the model output is structured, it is easy to automatically extract the YES/NO/MAYBE labels corresponding to each dimension of analysis and count them. If the predicted label is ‘YES’ and count(YES) + count (MAYBE) > count(NO), then the labels are consistent. Similarly, if the predicted label is ‘NO’ and count(NO) + count(MAYBE) > count(YES) then the labels are consistent.
We define $\textsc{Consistency}=\frac{\textsc{Cs-1}+\textsc{Cs-2}}{2}$. That is, a given output is consistent ($\textsc{Consistency}=1$) only if both Cs-1 and Cs-2 are $1$. We apply this metric to the silver rationales, and only retain those with a Consistency score of $1$.
2.3 Distillation into a smaller LM
We use the above filtered training to train/distill a smaller language model. Given a prompt $P$ (Cave’s prompt), a teacher model $T$ (GPT-4-Turbo) and a student model $S$ (LLaMa-3-8B), we generate silver training data:
@@ 2.3 Distillation into a smaller LM
where $x_{i}$ represents the input document pairs and $y_{i}^{t}$ represents the explanations created from $T$. We filter $D_{train}$ using the criteria of $\textsc{Accuracy}=1,\textsc{Consistency}=1$ and conformity to JSON/linguistic feature format to get $F_{train}$. We then train $S$ with $F_{train}$ with standard language modeling loss:
@@ 3.1 Datasets
We perform experiments on three datasets: IMDb62 Seroussi et al. (2014), Blog-Auth Schler et al. (2006) and FanFiction (via the PAN-2020 Authorship Verification task [^4] https://pan.webis.de/clef20/pan20-web/author-identification.html). For IMDb62 we use the test set of 1k text pairs created by Hung et al. (2023). For Blog-Auth, we create our own test set by sampling 1k document pairs from the original blog authorship corpus [^5] https://paperswithcode.com/dataset/blog-authorship-corpus. For FanFiction, we sample 1k document pairs from the PAN-2020 test set (which had 14k pairs).
@@
For IMDb62 and Blog-Auth, the original datasets consist of documents and their anonymous author IDs; we sample document pairs with equal amounts of same/different authors to create training datasets for them. For FanFiction, we take the ‘small’ training set provided by PAN-2020, and sample datapoints from the same. We start off with a training set of $\sim$1.8k document pairs; we then combine them with our prompt from Table 2 and get silver rationale data from GPT-4-Turbo (temperature $0.0$ and $2$ responses per sample). We then filter them according to the criteria in Section 2.2 and use them for training - our final training data sizes are 2.1k, 2.3k and 1.8k for IMDb62, Blog-Auth and FanFiction respectively.
@@ 3.2 Training and Evaluation details
We run experiments with LLaMa-3-8B Meta-AI (2024), which is orders of magnitude smaller than GPT-4-Turbo. We train it with Low-Rank Adaption (LoRA; Hu et al., 2021), a parameter-efficient training method which offers lightweight training no additional inference latency. We report all hyperparameters in Appendix C. We train models in both single dataset and multi-dataset settings and report metrics Accuracy on Consistency in Table 3. We also report out-of-domain metrics of single-dataset models in Section 3.6.
@@
PromptAV 73.0 - 62.2* - 57.0 -
Cave’s prompt 67.6 0.86 57.7 0.73 52.2 0.67
LLaMa-3-8B CoT Zero-shot 62.4 - 60.7 - 57.5 -
-Instruct PromptAV 62.8 - 57.9 - 53.8 -
Cave’s prompt 36.7 0.16 38.8 0.25 45.5 0.17
LLaMa-3-8B Cave Single-dataset 75.7* 0.97 60.4 0.94 60.2* 0.81
Multi-dataset 74.9 0.99* 57.9 0.96* 58.6 0.91*
@@ 3.3 Baselines
We present and compare with three strong zero-shot baselines. We conduct these experiments with both GPT-4-Turbo (temperature $0.0$) and LLaMa-3-8B-Instruct (greedy decoding).
  1. 1.
    Chain of Thought (CoT) ‘Let’s think step by step’ Kojima et al. (2022): We use the same prompt as in Hung et al. (2023), and we report the prompt in Appendix B. We measure Accuracy for this baseline by extracting the predicted confidence score666we do this by string matching for the presence of any decimal number in the output and thresholding it at $0.5$. Consistency cannot be calculated for this baseline.
  2. 2.
    PromptAV: We use the same prompt as in Hung et al. (2023) (also reported in Appendix B). The Accuracy and Consistency measurements are the same as for CoT.
  3. 3.
    Cave’s prompt: We use our prompt from Table 2. Accuracy is calculated based on the predicted ‘output label’, and Consistency is calculated as in Section 2.2.
Why show two zero-shot results?
In addition to showing the results of our baselines using a strong online model (GPT-4-Turbo), we also show zero-shot results using a less-strong local model (LLaMa-3-8B-Instruct). The performance degradation there shows that it is insufficient merely to take existing approaches and run them using a local model; actual changes to the approach (e.g. those we propose with Cave) are necessary to be able to run locally at a high level of accuracy and consistency.

### v2

@@ Abstract
Authorship Verification (Av) (do two documents have the same author?) is essential in many sensitive real-life applications. Av is often used in proprietary domains that require a private, offline model, making SOTA online models like ChatGPT undesirable. Current offline models however have lower downstream utility due to low accuracy/scalability (eg: traditional stylometry Av systems) and lack of accessible post-hoc explanations. In this work, we take the first step to address the above challenges with our trained, offline LLaMa-3-8B model Cave (Controllable Authorship Verification Explanations): Cave generates free-text Av explanations that are controlled to be (1) structured (can be decomposed into sub-explanations in terms of relevant linguistic features), and (2) easily verified for explanation-label consistency (via intermediate labels in sub-explanations). We first engineer a prompt that can generate silver training data from a SOTA teacher model in the desired Cave output format. We then filter and distill this data into a pretrained LLaMa-3-8B, our carefully selected student model. Results on three difficult Av datasets IMDb62, Blog-Auth, and FanFiction show that Cave generates high quality explanations (as measured by automatic and human evaluation) as well as competitive task accuracies111github.com/INK-USC/Controllable-AV-Explanations .
@@ 1 Introduction
Figure 1: Cave generates structured free-text explanations that can be automatically verified for consistency.
@@
Authorship Verification (Av) Koppel and Winter (2014) is the NLP task of determining if two input documents were written by the same author. Av is used for tasks such as plagiarism detection, forensic analysis (often in support of law enforcement), analysis of the spread of misinformation. Given the sensitive nature of this task, it is imperative to develop Av methods that are explainable, accessible and secure. Explainability ensures that the system’s decisions can be trusted by user. Accessibility ensures easier downstream usability; for example, a well-structured free-text explanation with clearly defined topic points is more easily understood than unstructured explanations such as long lists of low-level features. Finally, security is necessary to safely use the Av system for proprietary data; a locally-hosted model on the user’s server is more secure than an API call to an online model, which could potentially store the user’s sensitive data. Av is also an extremely difficult task for humans, in contrast to other NLP tasks such as machine translation; this further necessitates the need for automatic Av systems that accessible to the user, accurate and explainable.
@@
Early works performed Av by comparing hand-crafted features such as n-grams van Dam (2013), POS tags Moreau and Vogel (2013), LIWC features Uchendu et al. (2020); while these are explainable, they are hard to scale. Subsequent works employed neural architectures to improve scalability and accuracy, such as the usage of Siamese networks Araujo-Pino et al. (2020); Najafi and Tavan (2022); Boenninghoff et al. (2019) to assign authorship via distance between document embeddings. These methods, however, have little or no post-hoc explainability. This poses a huge problem: one cannot, for instance, make legal assertions about the authorship of a set of questionable documents solely on the basis of “a system said so” - decision makers (judges, university officials, intelligence analysts, etc.) need to know why.
@@
Recent years have seen the surge of large language models and their self-rationalization capabilities. Subsequently, there have been works Hung et al. (2023); Huang et al. (2024) that generate free-text rationales [^2] We use terms explanation and rationale interchangeably. for Av from models such as GPT-4 OpenAI (2024b) in a zero/few-shot manner. These methods however face issues such as unreliability of online models (varying results/details across runs, deprecation of APIs), inconsistency of rationales with predicted label Hung et al. (2023) (refer to Appendix B), and high expenses.
In this work, we propose our model Cave (Controllable Authorship Verification Explanations), an in-house model that generates authorship predictions along with high-quality, free-text rationales that explain the former post-hoc. The rationales are structured as a set of sub-explanations that reason about documents’ shared authorship in terms of Av-relevant linguistic features Hung et al. (2023); Boenninghoff et al. (2019); these sub-explanations have corresponding (intermediate) labels that provide further structure and also serve as a means to verify overall consistency with Cave’s authorship prediction.
Since there is no human-written Av explanation data that can be used to train our in-house model, we adopt the model distillation strategy: West et al. (2022); Li et al. (2023a, 2024) we engineer a prompt to generate rationales in the structure described above, generate silver-standard training data by prompting a large teacher model (GPT-4-Turbo OpenAI (2024a)), and use that to finetune a smaller student model (LLaMa-3-8B Meta-AI (2024)). We also introduce our automatic metric Consistency that measures the consistency between the rationale and the model’s final label; Consistency is used to filter the silver train data before finetuning Ramnath et al. (2023); Li et al. (2023a), and also assess Cave during test time. Our experiments with three difficult Av datasets show that Cave leads to competitive task performances as well as high quality of rationales.
@@ 2 Our Method
Our proposed pipeline consists of three parts: (1) generating silver rationales with a desired structure from GPT-4-Turbo via a carefully engineered prompt, ‘Cave’s prompt’, (2) filtering them according to our criteria of output format and Consistency, and (3) distilling these into a small language model to get the final Cave. Figure 2 shows a visual demonstration of the pipeline.
@@ 2.1 Cave’s prompt
To design Cave’s prompt, we started off by observing the strengths and weaknesses of prior prompting-based methods for Av explanations, PromptAV Hung et al. (2023), Lip Huang et al. (2024). The main strength of these prompts is their compilation of Av-relevant linguistic required for authorship analysis. The main weakness is the inaccessibility of their generated explanations towards automatic rationale quality evaluation333this is reinforced by these works reporting only task accuracy and no rationale quality metrics. Further, human analyses (Appendix B) of rationales generated by PromptAV and CoT (chain-of-thought Wei et al. (2022), a baseline self-rationalization method also presented in Hung et al. (2023)) show that these rationales can be potentially inconsistent with their final label, rendering them untrustworthy.
We adopted PromptAV as our base prompt, and performed several iterations of prompt engineering on GPT-4-Turbo alongside human analyses of rationale quality and downstream utility. Our final ‘Cave’s prompt’ (Table 2) generates a highly structured output rationale that can be easily decomposed for (1) verifying rationale-label consistency and (2) more accessible layman utility; Cave’s prompt retains all the linguistic features compiled by PromptAV Boenninghoff et al. (2019). As shown in Table 2, Cave’s outputs follow a JSON-type text format where the linguistic features are keys; for each feature, the model has to provide an explanation and an intermediate output (YES/NO/MAYBE) that denotes whether this feature indicates shared authorship. Finally, the model has to provide a score from 00 to $1$ which indicates its confidence that the documents have the same author, and a YES/NO task label output. The output’s JSON format enables us to easily parse the explanation into its different sub-steps in an automatic manner; (re-iterating) this enables the downstream utility and automatic evaluation of the rationale.
2.2 Automatic Evaluation & Filtering
We define the following two criteria to evaluate the structure and consistency of rationales. We use these criteria in the distillation process (Section 2.3) to filter out low-quality silver training data, and at evaluation time.
Structure.
We define the correct and complete output structure to be a JSON format with all the target linguistic features, final score and final label.
@@ 2.2 Automatic Evaluation & Filtering
Rationales that are inconsistent with the model’s predicted label are risky and unusable in downstream applications Lyu et al. (2023). Prior works in self-rationalization check for consistency either by human analysis Chen et al. (2023b) or by training models Hase et al. (2020); Chen et al. (2023a); Wiegreffe et al. (2021) that can do the same. In Cave’s output format, the presence of intermediate and final labels allows us to automatically verify rationale-label consistency. We define a Consistency metric using two binary $0/1$ variables Cs-1 and Cs-2:
Cs-1: Is the ‘final score’ (which indicates the model’s confidence in whether the two documents have the same author) consistent with the final label? If the final label is ‘YES’, is the final score $\geq 0.5$? Similarly, if the final label is ‘NO’, is the final score $\leq 0.5$?
Cs-2: Are the intermediate labels faithful to the final label? Since the model output is structured, it is easy to automatically extract the YES/NO/MAYBE labels corresponding to each dimension of analysis and count them. If the final label is ‘YES’ and count(YES) + count (MAYBE) > count(NO), then the labels are consistent. Similarly, if the final label is ‘NO’ and count(NO) + count(MAYBE) > count(YES) then the labels are consistent.
We define $\textsc{Consistency}=\frac{\textsc{Cs-1}+\textsc{Cs-2}}{2}$. That is, a given output is consistent ($\textsc{Consistency}=1$) only if both Cs-1 and Cs-2 are $1$.
2.3 Model Distillation
We use Cave’s prompt ($P$) to generate silver training data from GPT-4-Turbo ($T$); we then filter out the low-quality silver data based on output structure and Consistency. We distill this filtered training data into a smaller language model LLaMa-3-8B ($S$). Formally, given the prompt $P$, teacher model $T$ and student model $S$, we generate silver training data:
@@ 2.3 Model Distillation
where $x_{i}$ represents the input document pairs and $y_{i}^{t}$ represents the explanations created from $T$. We filter $D_{train}$ using the criteria of $\textsc{Consistency}=1$ and conformity to the complete JSON format to get $F_{train}$. We then train $S$ with $F_{train}$ with standard language modeling loss:
@@ 3.1 Datasets
We perform experiments on three popular Av datasets: IMDb62 Seroussi et al. (2014), Blog-Auth Schler et al. (2006) and FanFiction (via the PAN-2020 Authorship Verification task444pan-2020-task). Our datasets are diverse in terms of their domains (IMDb62- movie reviews, Blog-Auth- website blogs, FanFiction- movie/book-based fanfiction); we carefully select these three datasets in order to test the generalizability of our method.
For IMDb62 we use the test set of 1k text pairs created by Hung et al. (2023). For Blog-Auth, we create our own test set by sampling 1k document pairs from the original blog authorship corpus555blog-authorship-corpus. For FanFiction, we sample 1k document pairs from the PAN-2020 test set (which had 14k pairs).
@@
For IMDb62 and Blog-Auth, the original datasets consist of documents and their anonymous author IDs; we sample document pairs with equal amounts of same/different authors to create training datasets for them. For FanFiction, we take the ‘small’ training set provided by PAN-2020, and sample datapoints from the same. We start off with a training set of 1.8k document pairs per dataset. We then combine them with Cave’s prompt and get silver rationale data from GPT-4-Turbo (temperature $0.0$ and $2$ responses per sample); this results in an initial set of 3.6k datapoints. Finally, we filter them according to the criteria in Section 2.2 and use them for training - our final training data sizes are 2.3k, 2.7k and 2.1k for IMDb62, Blog-Auth and FanFiction respectively.
@@ 3.2 Training and Evaluation details
We distill the above silver training data into a LLaMa-3-8B, which is orders of magnitude smaller than GPT-4-Turbo. We train it with LoRA Hu et al. (2021), a parameter-efficient training method. We report all hyperparameters in Appendix C. We train models in both single dataset and multi-dataset settings and report metrics Accuracy, Consistency in Table 3.
@@
PromptAV Zero-shot 73.0 - 62.2 - 57.0 -
Cave’s prompt Zero-shot 67.6 0.86 57.7 0.73 52.2 0.67
Cave’s prompt Few-shot 71.3 0.94 64.0 0.97* 61.7* 0.86
CoT Zero-shot 62.4 - 60.7 - 57.5 -
LLaMa-3-8B PromptAV Zero-shot 62.8 - 57.9 - 53.8 -
-Instruct Cave’s prompt Zero-shot 36.7 0.16 38.8 0.25 45.5 0.17
Cave’s prompt Few-shot 53.8 0.98 62.3 0.99 48.1 0.93
LLaMa-3-8B CoT Single-dataset 62.0 - 64.2* - 61.5 -
Multi-dataset 75.3 - 62.5 - 57.6 -
PromptAV Single-dataset 63.8 - 58.4 - 58.2 -
Multi-dataset 68.6 - 62.7 - 63.3 -
LLaMa-3-8B Cave Single-dataset 75.7* 0.97 60.4 0.94 60.2 0.81
Multi-dataset 74.9 0.99* 57.9 0.96 58.6 0.91*
@@ 3.3 Baselines
We present and compare with multiple baselines:
Traditional stylometry.
We present a weak baseline Ikae (2021) from traditional stylometry. We represent input document pairs as the concatenation of their TF-IDF666sklearn.feature_extraction.text.TfidfVectorizer or word-count 777sklearn.feature_extraction.text.CountVectorizer features, and train a Support Vector Machine Cortes and Vapnik (1995) to classify their shared authorship status. We report final hyperparameters for the same in Appendix C.
Zero-shot, prior work.
We present three zero-shot baselines with GPT-4-Turbo (temperature $0.0$) and LLaMa-3-8B-Instruct (greedy decoding). The first baseline is chain-of-thought CoT Wei et al. (2022), a standard self-rationalization method that does not include any Av specific instructions. The second baseline is PromptAV Hung et al. (2023). We use the exact prompts for CoT and PromptAV as reported in Hung et al. (2023) (refer Appendix B). Both these baselines report confidence scores (0-1, where higher score means more confidence in shared authorship) in lieu of labels. We extract this confidence score by string matching for decimal numbers in the output; we measure Accuracy by thresholding these scores at $0.5$. Consistency cannot be calculated for these two baselines. The third zero-shot baseline is Cave’s prompt itself.
Finetuned.
Similar to how we distill Cave, we generate silver CoT and PromptAV data from GPT-4-Turbo, filter them according to task correctness and distill them into a LLaMa-3-8B. We report final train set sizes in Appendix C.
Few-shot.
Finally, we present a few-shot baseline with Cave’s prompt on GPT-4-Turbo and LLaMa-3-8B-Instruct. This baseline is strong, but undesirable owing to the high expenses of performing few-shot Av  [^8] Each individual document in our datasets has on-average 1000 tokens; performing even a 4-shot baseline is highly expensive.. We have submitted the few shot prompts in supplementary material.